OpenStudy (anonymous):

Can anyone help me prove the following trig identity? tan(50)-tan(40)=2tan(10) it is all in degrees.

OpenStudy (anonymous):

write tan(50) as tan(40+10) use formula tan(a+b)=tana+tanb/1-tanatanb

OpenStudy (anonymous):

tan 40+tan10/1-tan 40tan 10 -tan(40)=2tan(10)

OpenStudy (anonymous):

I am not getting the right answer for some reason.

OpenStudy (anonymous):

\[\cot (2\Pi/9) - \tan (40) = 2 \tan(10)\] \[\cot (2\Pi/9) = 2 \tan (10) + \tan (40)\] \[-2 \tan(10) - \tan (40) + \cot (2\pi/9)= 0\]

OpenStudy (anonymous):

@homeylova223 is that ur answer